OpenStudy (anonymous):

Justine earned $26,000 during her job at city hall. After a year, she recieved a 3% raise. Find her total earnings during the first five years on the job.

OpenStudy (richyw):

does she receive a 3% raise each year, or just after the first year?

OpenStudy (anonymous):

Just after the first year.

OpenStudy (richyw):

3% of $26000 is \(0.03\times26000\) so a 3% raise would be\[26000+0.03(26000)\]\[=26000(1+0.03)\]\[=1.03\times 26000\]

OpenStudy (richyw):

make sense?

OpenStudy (anonymous):

So, the answer would be 26,780?

OpenStudy (anonymous):

Lol

OpenStudy (richyw):

nope. That's what her salary will be per year after the raise.

OpenStudy (richyw):

she works 5 years, so you will need to add them all up. so if she works 4 years at the new salary and one at the old you need to add 4 of the new ones and one of the old. 4 times the old salary is\[4\times(1.03\times 26000)=4.12\times26000\]

OpenStudy (richyw):

so\[26000+4.12\times 26000\]\[=26000\times(1+4.12)\]\[=5.12\times26000\]

OpenStudy (richyw):

see what's going on?

OpenStudy (anonymous):

kind of

OpenStudy (richyw):

it would be a more interesting question if she got a raise each year.

OpenStudy (richyw):

ok if you make $5 a day for 5 days you make $25 right?

OpenStudy (richyw):

\(5+5+5+5+5=5\times 5=25\)

OpenStudy (richyw):

but if you make 5 dollars the first day and 6 dollars on the last day you make\(5+6+6+6+6=5+4\times6=29\)

OpenStudy (richyw):

that's all you are doing here. you just haven't been told how much she makes in the last 4 years, you have to calculate that from the percent.

OpenStudy (anonymous):

Oh! Okay! Thank you, I got it!